given that log10^7=x and log10^2=y, evaluate log10^35
Found 2 solutions by Theo, Edwin McCravy:
Answer by Theo(13342)

log(10^7) = x
log(10^5) = y

what is log 10^35 in terms of x and y?

log(10^7) = x if and only if 10^x = 10^7 which makes x = 7.

log(10^5) = y if and only if 10^y = 10^5 which makes y = 5.

log(10^35) = z if and only if 10^z = 10^35 which makes z = 35.

since z = 35, and 35 = 5 * 7, and y = 5 and x = 7, then you can say that z = x * y.

therefore, log(10^35) = x * y.

another way you can look at it is as follows:

log(10^7) = x
log(10^5) = y

log(10^35) is equal to log(10^(5*7) which is equal to log((10^5)^7) which is equal to 7 * log(10^5) which is equal to 7 * 5 * log(10) which is equal to 7 * 5 * 1 which is equal to 35.

since you know that x = 5 and y = 7, and you know that 35 = 5 * 7, then you know that this is equivalent to log(10^35) is equal to x * y.


Answer by Edwin McCravy(20077)



[You don't need to know anything about x or y or anything else.

Edwin